# Kinematic morphology of low-mass galaxies in IllustrisTNG

### v1

@@ Abstract
The origin of diverse kinematic morphologies observed in low-mass galaxies is unclear. In this study, we investigate the kinematic morphologies of central galaxies with stellar mass $10^{8.5-9.0}M_{\odot}$ at $z=0$ in the TNG50-1 cosmological simulation. The majority of the low-mass galaxies in TNG50-1 are dispersion-dominated, consistent with observations. By tracing the evolutionary histories of simulated low-mass galaxies, we find that while most stars form in rotating cold gas discs, the orientation of the star-forming discs relative to the galaxies may evolve with cosmic time. If the cold gas disc remains aligning with the galaxy during its evolution, stars formed at different times share the same rotational direction, leading to a rotation-dominated system. On the contrary, frequent misalignment of cold gas disc would result in a dispersion-dominated system. In addition, we also find that the two-body scattering can have a non-negligible numerical heating effect on the simulated galaxy morphology, especially at central regions of galaxies and for relatively low-mass galaxies. By comparing results of simulations with different resolutions, our results suggest that the simulated morphology of galaxies is reliable when their number of stellar particles exceeds about $10^{4}$, and bulge morphology of galaxies can not be resolved robustly at the resolution level of TNG50-1.
@@ 3.2 Misaligned Star Formation
In Fig. 5, we present $\mathrm{(B/T)}_{*}$ for all low-mass galaxies as a function of their mass-weighted misalignment angle ${\theta}^{{}^{\prime}}$. It can be seen that, there exists a clear correlation between the misalignment angle and the final stellar morphology of these low-mass galaxies. Specifically, low-mass galaxies with a larger ${\theta}^{{}^{\prime}}$, experiencing misaligned star formation more frequently, tend to become more dispersion-dominated with a larger $\mathrm{(B/T)}_{*}$ at the present day. And for those with a smaller ${\theta}^{{}^{\prime}}$, the newly-formed stars during evolution are usually aligned with the overall galaxy, resulting in a more rotation-dominated morphology. Therefore, Fig. 5 indicates that the extent of misalignment in star formation plays a crucial role in determining the final kinematic morphology of low-mass galaxies. Frequent misalignment of star formation would lead to a dispersion-dominated morphology, while a consistently aligned star formation history would in general form a disc-like low-mass galaxy.
@@ 4.2 Quantifying Two-Body Scattering Effect
The two-body scattering effect is commonly quantified by inspecting the two-body relaxation process theoretically (e.g. Binney & Tremaine, 1987, 2008), and also in simulations (e.g. Power et al., 2003; Diemand et al., 2004; Zhang et al., 2019; Ludlow et al., 2019a, b). In this work, we assess the two-body relaxation for simulated galaxies by computing their corresponding convergence radius, $r_{\mathrm{conv}}$. As discussed in Power et al. (2003) and Ludlow et al. (2019a, b), beyond $r_{\mathrm{conv}}$, the time required for two-body relaxation is comparable to or greater than the Hubble time, and the impact of two-body scattering is considered to be small. While within $r_{\mathrm{conv}}$, the two-body scattering effect can significantly affect the kinematics of simulated galaxies. To calculate the convergence radius, we basically follow the method of Ludlow et al. (2019a), except with further consideration of the contribution from stellar particles (see detailed calculation and explanation in Appendix A).
@@ 5 Conclusions and Discussions
The numerical effect that makes the simulated galaxies more dispersion-dominated than ideal cases affects more galaxies of lower masses, and galaxies simulated with lower resolutions. Therefore, cautions must be made when analysing the morphology of simulated galaxies with low-mass and/or low-resolution. While galaxies with half-mass radius larger than the convergence radius can be considered less affected by the two-body scattering effect, our results suggest that the simulated galaxy morphology becomes reliable when the number of stellar particles exceeds about $10^{4}$.
@@ Acknowledgements
We thank Aaron D. Ludlow and Shi Shao for their helpful discussions and comments. This work is supported by the National Natural Science Foundation of China (grant No. 11988101), the National SKA Program of China (Nos. 2022SKA0110200, 2022SKA0110201), the National Key Research and Development Program of China (No. 2023YFB3002500), the Strategic Priority Research Program of Chinese Academy of Sciences, (grant No. XDB0500203), and K.C. Wong Education Foundation. The IllustrisTNG simulations were undertaken with compute time awarded by the Gauss Centre for Supercomputing (GCS) under GCS Large-Scale Projects GCS-ILLU and GCS-DWAR on the GCS share of the supercomputer Hazel Hen at the High Performance Computing Center Stuttgart (HLRS), as well as on the machines of the Max Planck Computing and Data Facility (MPCDF) in Garching, Germany.

### v2

@@ Abstract
The origin of diverse kinematic morphologies observed in low-mass galaxies is unclear. In this study, we investigate the kinematic morphologies of central galaxies with stellar mass $10^{8.5-9.0}M_{\odot}$ at $z=0$ in the TNG50-1 cosmological simulation. The majority of the low-mass galaxies in TNG50-1 are dispersion-dominated, consistent with observations. By tracing the evolutionary histories of simulated low-mass galaxies, we find that while most stars form in rotating cold gas discs, the orientation of the star-forming discs relative to the galaxies may evolve with cosmic time. If the cold gas disc remains aligning with the galaxy during its evolution, stars formed at different times share the same rotational direction, leading to a rotation-dominated system. On the contrary, frequent misalignment of cold gas disc would result in a dispersion-dominated system. In addition, we also find that the two-body scattering can have a non-negligible numerical heating effect on the simulated galaxy morphology, especially at central regions of galaxies and for relatively low-mass galaxies. By comparing results of simulations with different resolutions, our results suggest that the simulated morphology of galaxies is roughly reliable when their number of stellar particles exceeds about $10^{4}$, and bulge morphology of galaxies can not be resolved robustly at the resolution level of TNG50-1.
@@ 3.2 Misaligned Star Formation
In Fig. 5, we present $\mathrm{(B/T)}_{*}$ for all low-mass galaxies as a function of their mass-weighted misalignment angle ${\theta}^{{}^{\prime}}$. It can be seen that, there exists a clear correlation between the misalignment angle and the final stellar morphology of these low-mass galaxies. Specifically, low-mass galaxies with a larger ${\theta}^{{}^{\prime}}$, experiencing misaligned star formation more frequently, tend to become more dispersion-dominated with a larger $\mathrm{(B/T)}_{*}$ at the present day. And for those with a smaller ${\theta}^{{}^{\prime}}$, the newly-formed stars during evolution are usually aligned with the overall galaxy, resulting in a more rotation-dominated morphology. Therefore, Fig. 5 indicates that the extent of misalignment in star formation plays a crucial role in determining the final kinematic morphology of low-mass galaxies. Frequent misalignment of star formation would lead to a dispersion-dominated morphology, while a consistently aligned star formation history would in general form a disc-like low-mass galaxy. Similar mechanisms have been found in more massive galaxies in previous studies of numerical simulations (e.g. Sales et al., 2012; Park et al., 2019, 2022; Casanueva et al., 2022).
@@ 4.2 Quantifying Two-Body Scattering Effect
The two-body scattering effect is commonly quantified by inspecting the two-body relaxation process theoretically (e.g. Binney & Tremaine, 1987, 2008), and also in simulations (e.g. Power et al., 2003; Diemand et al., 2004; Zhang et al., 2019; Ludlow et al., 2019a, b). In this work, we assess the two-body relaxation for simulated galaxies by computing their corresponding convergence radius, $r_{\mathrm{conv}}$. As discussed in Power et al. (2003) and Ludlow et al. (2019a, b), beyond $r_{\mathrm{conv}}$, the time required for two-body relaxation is comparable to or greater than the Hubble time, and the impact of two-body scattering is considered to be small. While within $r_{\mathrm{conv}}$, the two-body scattering effect can significantly affect the kinematics of simulated galaxies.
To calculate the convergence radius, we basically follow the method of Ludlow et al. (2019a), except with further consideration of the contribution from stellar particles. For simulated galaxies, the two-body relaxation time $t_{\mathrm{relax}}$ is defined as the time required for significant velocity change of a particle due to cumulative two-body encounters with other particles. Based on it, the convergence radius $r_{\mathrm{conv}}$ of simulated galaxies is estimated by requiring the two-body relaxation time at the convergence radius $t_{\mathrm{relax},r_{\mathrm{conv}}}$ to be equal to the orbital time at the virial radius $t_{\mathrm{orbit},200}$, i.e., $t_{\mathrm{relax},r_{\mathrm{conv}}}/t_{\mathrm{orbit},200}=1$. More details on the calculations can be found in Appendix A.
@@ 5 Conclusions and Discussions
The numerical effect that makes the simulated galaxies more dispersion-dominated than ideal cases affects more galaxies of lower masses, and galaxies simulated with lower resolutions. Therefore, cautions must be made when analysing the morphology of simulated galaxies with low-mass and/or low-resolution. While galaxies with half-mass radius larger than the convergence radius can be considered less affected by the two-body scattering effect, our results suggest that the simulated galaxy morphology becomes roughly reliable when the number of stellar particles exceeds about $10^{4}$.
@@ Acknowledgements
We thank the anonymous referee for the prompt and constructive report. We thank Aaron D. Ludlow and Shi Shao for their helpful discussions and comments. This work is supported by the National Natural Science Foundation of China (grant No. 11988101), the National SKA Program of China (Nos. 2022SKA0110200, 2022SKA0110201), the National Key Research and Development Program of China (No. 2023YFB3002500), the Strategic Priority Research Program of Chinese Academy of Sciences, (grant No. XDB0500203), and K.C. Wong Education Foundation. The IllustrisTNG simulations were undertaken with compute time awarded by the Gauss Centre for Supercomputing (GCS) under GCS Large-Scale Projects GCS-ILLU and GCS-DWAR on the GCS share of the supercomputer Hazel Hen at the High Performance Computing Center Stuttgart (HLRS), as well as on the machines of the Max Planck Computing and Data Facility (MPCDF) in Garching, Germany.